OpenStudy (anonymous):

You have £1,000 in your savings account. At the end of each month, the bank adds an extra 0.2% of interest to the outstanding balance and you make a withdrawal of £10. How much is in the savings account after this has happened for 36 months, rounded to the nearest pound? (I guess you need to solve it with a linear difference equation, but I cannot figure it out)

OpenStudy (anonymous):

\[a ^{n}x _{0} +b (a ^{n}-1)/(a-1)\] I found this formula for it and it gives the good answer but I do not see why. (a^n=1.002^36 , xo=1000 , b=-10)

OpenStudy (anonymous):

yeah i have to write this down. certainly cannot do it on the fly. if at all!

OpenStudy (anonymous):

wait, why isn't it just \[1000(1.002)^{10}-10(1.002)^9\]?

OpenStudy (anonymous):

actually my exponents are ridiculous. it is 36 months not ten. so i think maybe \[1000(1.002)^{36}-10(1.002)^{35}\]

OpenStudy (anonymous):

i just wrote out the first three months and saw what i got.

OpenStudy (anonymous):

I am not sure here at all, I found that formula and it did give me the correct answer, but I can only see the answer unfortunately

OpenStudy (anonymous):

first month P(1.002) second month (P(1.002)-10)(1.002)=P(1.002)^2-10(1.002) third month (P(1.002)^2-10(1.002))-10)(1.002)=P(1.002)^3-10(1.002)^2-10(1.002)

OpenStudy (anonymous):

oh wait maybe you get a geometric series for the ten part.

OpenStudy (anonymous):

yeah that formula does look like a geometric series

OpenStudy (anonymous):

first part is definitely p(1.002)^36

OpenStudy (anonymous):

lets see what the last one is.

OpenStudy (anonymous):

I think I see now why it is so

OpenStudy (anonymous):

guess it depends on whether you take out ten pounds at the end. if so you will get 10+10(1.002)+10(1.002)^2+...+10(1.002)^35

OpenStudy (anonymous):

\[x _{n}=a ^{n} x _{0} +(a ^{n-1} +....+a+1)b\]

OpenStudy (anonymous):

and the last part is a series

OpenStudy (anonymous):

equal to a^n -1/a-1

OpenStudy (anonymous):

yeah i think so. i bet this simplifies to get essentially what i wrote at the beginning. if my algebra was good enuf i could do it

OpenStudy (anonymous):

did you try 1000(1.002)^36-10(1.002)^35?

OpenStudy (anonymous):

no, but I can

OpenStudy (anonymous):

that gives a positive amount, that cannot be right, the interest is around 2 and you take out 10

OpenStudy (anonymous):

we can use 1+r+r^2+...+r^35=\[\frac{r^{36}-1}{1-r}\] with r = 1.002

OpenStudy (anonymous):

3.7289 rounded

OpenStudy (anonymous):

yes that is what I thought too, but isnt the denominator r-1 ?

OpenStudy (anonymous):

multiply by ten to get 37.289

OpenStudy (anonymous):

yea typo

OpenStudy (anonymous):

denominator in this case is .002

OpenStudy (anonymous):

oh wait my calculation is wrong

OpenStudy (anonymous):

i divided by .02

OpenStudy (anonymous):

the answer was something around 700

OpenStudy (anonymous):

i am off by a decimal

OpenStudy (anonymous):

:) happens

OpenStudy (anonymous):

\[\frac{(1.002)^{36}-1}{.002}=37.289\]

OpenStudy (anonymous):

according to my calculator.

OpenStudy (anonymous):

multiply by ten to get 372.89

OpenStudy (anonymous):

according to my brain

OpenStudy (anonymous):

subtract from 1000(1.002)^36 = 1074.578 rounded

OpenStudy (anonymous):

gives the correct answer! sweet

OpenStudy (anonymous):

and i get as a final answer (as they say) 701.68

OpenStudy (anonymous):

whew.

OpenStudy (anonymous):

glad we got this worked out even if a couple of false starts

OpenStudy (anonymous):

maths is like this, try hard and fail a lot